Calcium hydride, CaH2, reacts with water to form hydrogen gas.

CaH2(s) + 2 H2O(l) Ca(OH)2(aq) + 2 H2(g)

This reaction is sometimes used to inflate life rafts, weather balloons, and the like, where a simple, compact means of generating H2 is desired. How many grams of CaH2 are needed to generate 14.0 L of H2 gas if the pressure of H2 is 730. torr at 27°C? g
Answer #1

fos Ha gas v= 142, T= 27C = 273+27 = 300k (1 atm= 760 tosor) p = 730 tors = 130 760 0.g6 atm = 0.082 tam mo0-х pU=ART , R = 0